OpenStudy (anonymous):

Need help? ._. Simplify the expression: 14 + 8 • 2. A. 30 B. 42 C. 11 D. 36

OpenStudy (anonymous):

A

OpenStudy (anonymous):

PEMDAS. 8x2 = 16 16+14 = 30

OpenStudy (anonymous):

I need help on some more please help

OpenStudy (anonymous):

@ferrariwindows77

OpenStudy (anonymous):

Once again. Do the parenthesis first. Multiplication second, then sub/add.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

no A

OpenStudy (anonymous):

yes its A

OpenStudy (anonymous):

help me on 2 more I'll fan both of you :)

OpenStudy (anonymous):

@ferrariwindows77

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A once again

OpenStudy (anonymous):

Write a variable expression that represents the word problem. Matt gave q quarters and n nickels to his sister. Which expression represents the amount in dollars that Matt gave to his sister? A. (0.25 + 0.05) • qn B. 0.25q + 0.05n C. q • n D. q + n

OpenStudy (anonymous):

B

OpenStudy (anonymous):

no im sorry D

OpenStudy (anonymous):

3 more

OpenStudy (anonymous):

Which expression completes this equation? 20 ÷ 4 + 2 • 8 = ? A. 16 – [3 + (4 – 1)] B. 18 • 2 – 10 ÷ 5 C. 20 ÷ 6 • 8 D. [(6 – 2) + 3] • 3

OpenStudy (anonymous):

@Starr_DynastyT

OpenStudy (anonymous):

@ferrariwindows77

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Which value of x makes the open sentence true? 5x – 6 = x + 12 A. 1 B. 1.5 C. 3 D. 4.5

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Which equation represents this situation? Naila worked 3.5 hours more than Lisa. Together, they worked a total of 18 hours. Lisa worked h hours. A. h + 3.5 = 18 B. h + 7 = 18 C. h + (h – 3.5) = 18 D. h + (h + 3.5) = 18

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Andrew wants to place grouping symbols in the expression 15 + 25 – 18 ÷ 2 so that it has a value of 18.5. Which shows the correct placing of the grouping symbols? A. (15 + 25) – 18 ÷ 2 B. 15 + (25 – 18 ÷ 2) C. 15 + (25 – 18) ÷ 2 D. 15 + 25 – (18 ÷ 2)

OpenStudy (anonymous):

@Starr_DynastyT